# Fast nonparametric inference of network backbones for graph sparsification

### v1

@@ II.3 Bayesian Generalization
$\times$ $B(k_{i}^{(b)}+1,s_{i}^{(b)}-k_{i}^{(b)}+1)$
$\times$ $B(k_{i}-k_{i}^{(b)}+1,(s_{i}-s_{i}^{(b)})-(k_{i}-k_{i}^{(b)})+1),$
where $P(\partial_{i},\{b_{ij}\}_{(i,j)\in\partial_{i}})$ is the local marginal joint likelihood of the neighborhood $\partial_{i}$ and its backbone edges $\{b_{ij}\}_{(i,j)\in\partial_{i}}=\partial^{(b)}_{i}$ after integrating over $\{\pi_{b}(i),\theta_{1}(i),\theta_{0}(i)\}$ with a uniform prior. (As before, the local model can be directly obtained from the global model by mapping the variables $\{W,E,W^{(b)},E^{(b)}\}\to\{s_{i},k_{i},s_{i}^{(b)},k_{i}^{(b)}\}$ and treating each neighborhood $\partial_{i}$ separately.) We can then a similar substitution as above to find
$-\log P^{(\text{local})}(G,\bm{b})$ $=\mathcal{L}^{(\text{local})}(G^{(b)})$ (14)
$+O\Big{(}\sum_{i=1}^{N}[\log s_{i}+\log k_{i}]\Big{)},$
with $\mathcal{L}^{(\text{local})}(G^{(b)})$ given by Eq. 5 and the Big-O notation used similarly as above.
Matching terms with Eq. II.2, we can see that the neighborhood-level description length exceeds $\log s_{i}+\log k_{i}$ by at least $\log{s_{i}-1\choose k_{i}-1}$, and so in the limit of $s_{i},k_{i}\gg 1$ the fractional error between $-\log P^{(\text{local})}(G,\bm{b})$ and $\mathcal{L}^{(\text{local})}(G^{(b)})$ becomes vanishingly small. Therefore MAP estimation with this Bayesian model is asymptotically equivalent to the MDL minimization problem with the objective in Eq. 5.
Generalizing beyond geometric weight distributions, we can formulate a flexible fully Bayesian generative model for global network backbones as follows, starting with an unweighted network:
@@
    Draw an inclusion probability $\pi_{b}$ according to a Beta prior $P(\pi_{b})$, which may be uniform. Assign each edge $(i,j)$ to the backbone $G^{(b)}$ with probability $\pi_{b}$ to construct the indicator vector $\bm{b}$.
@@
    Choose a discrete weight distribution $P(w_{ij}|\theta_{b_{ij}})$. Draw $\theta_{0},\theta_{1}$ from a factorizable prior $P(\theta_{0},\theta_{1})=P(\theta_{0})P(\theta_{1})$ that is conjugate to $P(w_{ij}|\theta_{b_{ij}})$ and may be uniform. Set the $\theta$ that gives a higher mean weight as $\theta_{1}$ if generating a high weight backbone.
  3. 3.
    Generate the weight $w_{ij}$ for each existing edge $(i,j)$ according to $P(w_{ij}|\theta_{b_{ij}})$.
We can also formulate a Bayesian generative model for local network backbones as follows:
  1. 1.
    For each neighborhood $\partial_{i}$, draw an inclusion probability $\pi_{b}(i)$ according to a Beta prior $P(\pi_{b}(i))$, which may be uniform. Assign each edge $(i,j)\in\partial_{i}$ to the backbone $G^{(b)}$ with probability $\pi_{b}(i)$ to construct the indicator vector $\bm{b}$.
  2. 2.
    Choose a discrete weight distribution $P(w_{ij}|\theta_{b_{ij}}(i))$ for each neighborhood $\partial_{i}$. Draw $\theta_{0}(i),\theta_{1}(i)$ from a factorizable prior $P(\theta_{0}(i),\theta_{1}(i))=P(\theta_{0}(i))P(\theta_{1}(i))$ that is conjugate to $P(w_{ij}|\theta_{b_{ij}}(i))$. The distributions $P(w_{ij}|\theta_{b_{ij}}(i))$ may be different for different nodes if desired.
  3. 3.
    Generate the weight $w_{ij}$ for each existing edge $(i,j)$ according to $P(w_{ij}|\theta_{b_{ij}}(i))$.
Both of these frameworks permit nonparametric MAP estimation by minimizing the description length $-\log P(G,\bm{b})$ over labellings (backbones) $\bm{b}$. The joint marginal likelihood $P(G,\bm{b})$ can be computed by integrating out the model parameters which have the appropriate conjugate priors for easy marginalization. This general framework also allows for model selection among both weight distributions and global or local schemes to determine the optimal backbone for a given network. However, depending on the weight distribution one chooses, the optimization may not be as simple as the greedy scheme described here. In these cases one can minimize over $\bm{b}$ using discrete MCMC methods or perhaps with continuous optimization methods for a relaxation of $\bm{b}$ to a real-valued vector.

### v3

@@ II.1 Canonical Models for Global and Local Backboning
×$\times\prod_{i=1}^{N}\int P(\bm{\theta}_{1}(i))\prod_{e\in\partial^{(b)}_{i}}P(w_{e}|\bm{\theta}_{1}(i))d\bm{\theta}_{1}(i)$
×$\times\prod_{i=1}^{N}\int P(\bm{\theta}_{0}(i))\!\!\prod_{e\in\partial_{i}\setminus\partial^{(b)}_{i}}\!\!P(w_{e}|\bm{\theta}_{0}(i))d\bm{\theta}_{0}(i),$
where $\bm{b}_{\partial_{i}}=\{b_{e}:e\in\partial_{i}\}$ is the backbone partition of the edges in the neighborhood $\partial_{i}$. The description length of interest is then
$\mathcal{L}^{(\text{local})}_{C}(G^{(b)})=-\log P^{(\text{local})}_{C}(G,\bm{b})=\sum_{i=1}^{N}\mathcal{L}_{C}^{(\text{neig})}(\partial_{i}^{(b)}),$ (10)
where
$\mathcal{L}^{(\text{neig})}_{C}$ $(\partial_{i}^{(b)})=\log(k_{i}+1)+\log{k_{i}\choose k_{i}^{(b)}}$
$~{}~{}-\log\left[\int P(\bm{\theta}_{1}(i))\prod_{e\in\partial^{(b)}_{i}}P(w_{e}|\bm{\theta}_{1}(i))d\bm{\theta}_{1}(i)\right]$ (11)
$~{}~{}-\log\left[\int P(\bm{\theta}_{0}(i))\!\!\prod_{e\in\partial_{i}\setminus\partial^{(b)}_{i}}\!\!P(w_{e}|\bm{\theta}_{0}(i))d\bm{\theta}_{0}(i)\right].$
Similarly to the global case, the local objective has bit flip symmetry, but in this case at the neighborhood level. We therefore focus on the regime $k_{i}^{(b)}\leq k_{i}/2$, to promote sparse backbones—under the same logic as before, one cannot say that more than half of the edges in a neighborhood have statistically significant weights relative to the rest of the neighborhood, under any reasonable null model that treats nodes independently.
Eq. II.1 and Eq. 10 provide flexible objectives to solve for the backbone $G^{(b)}$ of a weighted network $G$ by assessing global and local edge weight heterogeneity respectively. These canonical model objectives allow for any underlying assumption about the weight distributions within and outside of the backbone $G^{(b)}$. They are adaptable to either discrete or continuous weights, and provide a means of model selection among different mixtures of underlying weights through the comparison of the description length values under different distributional assumptions for $P$.
II.2 Microcanonical Models for Global and Local Backboning
An alternative Bayesian formulation of network backbones can be done in a microcanonical framework in which we fix the network weight $W$ rather than observing $W$ as a result of generating the network edge weights. This allows us to calibrate the model so that the weight generative process being fit to the observed data $G$ will always generate the correct total weight $W$. This has the benefit of not requiring the marginalization over any free parameters. Additionally, fixing statistics of interest exactly rather than in expectation (and vice versa) can permit easier analytical treatment as well as result in qualitatively distinct statistical behaviors in network ensembles voitalov2020weighted , making microcanonical models worth exploring in their own right.
For the microcanonical generative model of network backbones, we assume that $W$ is known in addition to the unweighted graph $G_{u}$. We first draw the number of backbone edges $E^{(b)}$ uniformly at random from the interval $[0,E]$, giving a prior of $P(E^{(b)})=(E+1)^{-1}$. Similarly, we draw the total backbone weight $W^{(b)}$ uniformly at random. Given that we have positive integer weights, we know that $W^{(b)}\geq E^{(b)}$, and $W^{(b)}\leq W-(E-E^{(b)})$, since each edge must have weight at least $1$. Therefore, given $E^{(b)}$, $W^{(b)}$ is then drawn uniformly at random from the interval $[E^{(b)},W-(E-E^{(b)})]$, resulting in the prior $P(W^{(b)}|E^{(b)})=(W-E+1)^{-1}$. We then draw the backbone assignments $\bm{b}$ uniformly at random from the set of all binary partitions with $\sum_{e=1}^{E}b_{e}=E^{(b)}$, of which there are (${E\choose E^{(b)}}$, giving a prior of $P(\bm{b}|E^{(b)})={E\choose E^{(b)}}^{-1}$. Finally, given the specification of the backbone edges and total backbone weight, we draw the weights of all edges in the backbone and non-backbone uniformly at random given that the $E^{(b)}$ edges in the backbone must have weights that sum to $W^{(b)}$ and the $E-E^{(b)}$ edges in the non-backbone must have weights that sum to $W-W^{(b)}$. The number of configurations of $k$ positive integers that sum to $n$ is called the number of compositions of the integer $n$ into $k$ parts, and is given by (${n-1\choose k-1}$, so the probability of a particular configuration of the backbone and non-backbone weights given their sums is $P(G|\bm{b},W^{(b)},E^{(b)})={W^{(b)}-1\choose E^{(b)}-1}^{-1}{W-W^{(b)}-1\choose E-E^{(b)}-1}^{-1}$.
Putting this all together, the full joint distribution of the microcanonical model can be written as
$P^{(\text{global})}_{M}(G,\!\bm{b})\!$ $=\!P(E^{(b)})\!\times\!P(W^{(b)}|E^{(b)})\!\times\!P(\bm{b}|E^{(b)})$
×$~{}~{}~{}~{}~{}\!\times\!P(G|\bm{b},W^{(b)},E^{(b)})$ (12)
$\!=\!(E\!+\!1)^{-1}\!\times\!(W\!-\!E\!+\!1)^{-1}\!\times\!{E\choose E^{(b)}}^{-1}$
×($~{}~{}~{}~{}~{}\!\times\!{W^{(b)}\!-\!1\choose E^{(b)}\!-\!1}^{-1}{W\!-\!W^{(b)}\!-\!1\choose E\!-\!E^{(b)}\!-\!1}^{-1},$
which has description length
$\mathcal{L}_{M}^{(\text{global})}\!(G^{(b)})\!$ $=\!-\!\log P^{(\text{global})}_{M}(G,\bm{b})$ (13)
$\!=\!\log(E+1)\!+\!\log(W\!-\!E\!+\!1)\!+\!\log{E\choose E^{(b)}}$
$~{}+\log{W^{(b)}\!-\!1\choose E^{(b)}\!-\!1}\!+\!\log{W-W^{(b)}\!-\!1\choose E\!-\!E^{(b)}-1}.$
As before, this objective can be minimized over backbones $G^{(b)}$ to find the MDL-optimal backbone under the microcanonical global model.
We can also formulate a microcanonical local backbone model in a similar way, drawing $\{k_{i}^{(b)},s_{i}^{(b)},\bm{b}_{\partial_{i}},\partial^{(b)}_{i},\partial_{i}\setminus\partial^{(b)}_{i}\}$ using hierarchical uniform priors analogous to those for $\{E^{(b)},W^{(b)},\bm{b},G^{(b)},G\setminus G^{(b)}\}$ respectively, to form each neighborhood-level term $P^{(\text{global})}_{M}(\partial_{i}^{(b)},\bm{b}_{\partial_{i}}|s_{i})$ in an independent factorization analogous to Eq. II.1. The only additional consideration we must make for the microcanonical model is that, since $s_{i}$ is assumed known for all $i=1,...,N$, we need to put a uniform prior on $\bm{s}=\{s_{1},...,s_{N}\}$ given the known total network weight $W$. By adding this prior, we can compare the global and local microcanonical description lengths directly, since both models assume the same prior knowledge ($G_{u}$ and $W$). Since each edge has weight at least $1$, node $i$ must have a strength that satisfies $s_{i}\geq k_{i}$. In other words, a uniform distribution of strengths $\bm{s}$ will consider all possible ways to distribute the excess weight $W-E$ onto the $N$ nodes, allowing for some nodes to receive no excess weight (in other words, $s_{i}=k_{i}$). A uniform prior on $\bm{s}$ is then equivalent to generating a random composition of $W-E$ as $N$ non-negative integers, allowing the integers to potentially be zero, of which there are (${N+W-E-1\choose W-E}$ possible combinations. The final prior for $\bm{s}$ is then $P(\bm{s})={N+W-E-1\choose W-E}^{-1}$. This prior is a global constant that does not impact the inferred backbones, but is technically necessary for a direct comparison of the global and local description length values for model comparison.
Including the prior on the strengths $P(\bm{s})$, we have a description length for the local microcanonical model of
$\mathcal{L}^{(\text{local})}_{M}$ $(G^{(b)})=\!-\!\log P(\bm{s})\!-\!\log P^{(\text{local})}_{M}(G^{(b)},\bm{b}|\bm{s})$ (14)
$=\!-\!\log P(\bm{s})\!-\!\log\left(\prod_{i=1}^{N}P^{(\text{global})}_{M}(\partial_{i}^{(b)},\!\bm{b}_{\partial_{i}}|s_{i})\right)$
$=\log{N\!+\!W\!-\!E\!-\!1\choose W\!-\!E}+\sum_{i=1}^{N}\mathcal{L}_{M}^{(\text{neig})}(\partial_{i}^{(b)}),$
where
$\mathcal{L}^{(\text{neig})}_{M}(\partial_{i}^{(b)})$ $=\log(k_{i}\!+\!1)+\log(s_{i}\!-\!k_{i}\!+\!1)+\log{k_{i}\choose k_{i}^{(b)}}$
$~{}+\log{s_{i}^{(b)}-1\choose k_{i}^{(b)}-1}+\log{s_{i}-s_{i}^{(b)}-1\choose k_{i}-k_{i}^{(b)}-1}.$ (15)
One can show (see Appendix A) that the microcanonical global and neighborhood-level description lengths of Eq. 13 and Eq. II.2 are asymptotically equivalent to the canonical description lengths of Eq. II.1 and Eq. II.1 respectively when weights follow a geometric distribution. The full canonical local description length of Eq. 10 with geometric weights is only equivalent to the full microcanonical local description length of Eq. 14 up to the term $\log{N+W-E-1\choose W-E}$ associated with the prior $P(\bm{s})$, which is a global constant that does not affect the inferred backbones.
One of the major benefits of using MDL-based learning objectives is that they naturally provide a simple criterion for model selection—the model providing the lowest description length at the end of the fitting process is the best model for the data rissanen1978modeling . We can then determine whether a global or local backbone, as well as which weight model, provides the most parsimonious description of the observed network $G$, by fitting the objectives in Eq. II.1, Eq. 10, Eq. 13, and Eq. 14 then comparing the resulting description lengths. The common downside of MDL objectives is that they can be challenging to minimize, often requiring sophisticated Markov chain Monte Carlo methods peixoto2017nonparametric . However, we show in the next section that the microcanonical description lengths in Eq. 13 and Eq. 14 can be minimized exactly using a fast greedy algorithm. We can extend this proof of optimality to a set of common weight distributions under the canonical formulations in Eq. II.1 and Eq. 10.
II.3 Optimization
All four of the MDL backboning objectives of Sec. II.1 and Sec. II.2—Eq.s II.11013, and 14—must be minimized over sets of backbone edges $G^{(b)}\subseteq G$ to identify an MDL-optimal backbone under each description length formulation. With naïve optimization by enumeration of all possible backbones $G^{(b)}$, this would require $2^{E}$ evaluations of a given objective—one for each possible backbone—which is intractable for all but the smallest networks. We show here that for the microcanonical objectives of Eq.s 13 and 14, as well as for a wide variety of weight distributions for the canonical models of Eq.s II.1 and 10, one can find the MDL-optimal backbone using simple greedy procedures requiring only $O(E\log E)$ operations—the bottleneck consisting of sorting the set of edges $G$ by weight at the beginning of the algorithm.
We can analyze the microcanonical objectives first, then generalize our results to the canonical objectives. Consider a greedy procedure in which the backbone $G^{(b)}$ starts empty and we add edges $e\in G$ to $G^{(b)}$ in decreasing order of weight $w_{e}$, stopping the procedure when $E^{(b)}=E/2$ edges have been added to the backbone. (As discussed in Sec. II.1, given the bit flip symmetry of the description length objectives with respect to $\bm{b}$, the regime of interest for backboning is $E^{(b)}\leq E/2$.) We then scan over the past solution candidates and identify the one with the lowest description length to find the optimal backbone. Details of this procedure are shown in Algorithm 1. During this algorithm, at all steps we must necessarily have that
$\frac{W^{(b)}}{E^{(b)}}\geq\frac{W}{E}\geq\frac{W-W^{(b)}}{E-E^{(b)}}.$ (16)
In other words, the average edge weight in the backbone $G^{(b)}$ is at least as high as the average edge weight in the total network $G$, which is in turn at least as high as the average edge weight in the non-backbone $G\setminus G^{(b)}$.
We can now consider the change in description length
$\Delta\mathcal{L}_{M}$ ${}^{(\text{global})}(W^{(b)}\mathrel{+}=1)=\mathcal{L}_{M}^{(\text{global})}(G_{+}^{(b)})-\mathcal{L}_{M}^{(\text{global})}(G^{(b)})$
$=\log{\frac{W^{(b)}}{W^{(b)}-E^{(b)}+1}\frac{(W-W^{(b)})-(E-E^{(b)})}{W-W^{(b)}-1}},$ (17)
where $E_{+}^{(b)}=E^{(b)}\geq 1$ and $W_{+}^{(b)}=W^{(b)}+1$ for the backbone $G_{+}^{(b)}$. Eq. II.3 gives the change in description length induced by an increase in the total backbone weight $W^{(b)}$ for a fixed number of backbone edges $E^{(b)}$. Rearranging the argument of the logarithm, we can see that $\Delta\mathcal{L}_{M}^{(\text{global})}(W^{(b)}\mathrel{+}=1)<0$—in other words, the description length is decreasing in $W^{(b)}$—when
$W^{(b)}>(E^{(b)}-1)\frac{W-1}{E-2}.$ (18)
Now, using the inequality of Eq. 16 and $21\leq E^{(b)}\leq E/$ 2, we have that for the greedy procedure in Algorithm 1
$W^{(b)}$ $=(E^{(b)}-1)\frac{W^{(b)}}{E^{(b)}}\frac{E^{(b)}}{E^{(b)}-1}$
$\geq(E^{(b)}-1)\frac{W}{E}\frac{E}{E-2}$
$>(E^{(b)}-1)\frac{W-1}{E-2},$ (19)
which is precisely the condition required in Eq. 18 for the description length $\mathcal{L}_{M}^{(\text{global})}$ to be deceasing in the backbone weight $W^{(b)}$ for fixed $E^{(b)}$. In other words, the best backbone $G^{(b)}$ for any number of edges $E^{(b)}\in[1,E/2]$ is the one with the highest total weight $W^{(b)}$, and so Algorithm 1 will always be able to find an optimal solution for the backbone $G^{(b)}$. (The other optimal solution has backbone $G\setminus G^{(b)}$, due to the bit flip symmetry.) Note that this condition still allows for the backbone $G^{(b)}$ to potentially be empty—the description length may be minimized at $E^{(b)}=0$, if the weights are sufficiently homogeneous. Using a similar argument, we can also show that the canonical description length of Eq. II.1 is minimized with the greedy procedure in Algorithm 1 for common weight distributions including the Poisson, Geometric, and Exponential distributions, which all fall under the Natural Exponential Family (NEF) of distributions morris1982natural (see Appendix D).
Mapping the global variables $\{E^{(b)},W^{(b)},\bm{b},G^{(b)},G\setminus G^{(b)}\}$ to the neighborhood level variables $\{k_{i}^{(b)},s_{i}^{(b)},\bm{b}_{\partial_{i}},\partial^{(b)}_{i},\partial_{i}\setminus\partial^{(b)}_{i}\}$ as before, we have that the best neighborhood-level backbone $\partial_{i}^{(b)}$ for any node degree $k_{i}^{(b)}\in[1,k_{i}/2]$ is the one with the highest total weight $s_{i}^{(b)}$, and so the application of the greedy procedure in Algorithm 1 to each neighborhood $\partial_{i}$ separately gives the MDL-optimal backbone $G^{(b)}=\cup_{i=1}^{N}\partial_{i}^{(b)}$ for the local objectives of Eq. 10 and Eq. 14. Details of this procedure are shown in Algorithm 2.
Input: Graph $G=\{(e=(i,j),w_{e}=w_{ij})\}_{e=1}^{E}$
Output: Global backbone $G^{(b)}\subseteq G$
Initialization:
-Sort edges $G$ by decreasing edge weight $w_{e}$;
Algorithm:
-Set $W^{(b)}=0$, $G^{(b)}=[~{}]$, $\mathcal{L}s=[~{}]$;
-Compute $\mathcal{L}=\mathcal{L}^{(\text{global})}(W^{(b)}=0,E^{(b)}=0)$ using Eq. II.1 or Eq. 13 (whichever is being minimized) and append to $\mathcal{L}s$;
for $E^{(b)}\leftarrow 1$ to $\lfloor E/2\rfloor$ do
             $i,j,w_{ij}\leftarrow G[E^{(b)}-1]$;
             $W^{(b)}\mathrel{+}=w_{ij}$;
             Append $(i,j,w_{ij})$ to $G^{(b)}$;
             $\mathcal{L}\leftarrow\mathcal{L}^{(\text{global})}(W^{(b)},E^{(b)})$;
             Append $\mathcal{L}$ to $\mathcal{L}s$;
end for
-Set $G^{(b)}\leftarrow G^{(b)}[:\text{argmin}(\mathcal{L}s)]$;
return $G^{(b)}$;
Algorithm 1 Global backboning (Eq. II.1, Eq. 13)
Input: -Graph $G=\{(e=(i,j),w_{e}=w_{ij})\}_{e=1}^{E}$;
            -directed (boolean)
Output: Local backbone $G^{(b)}\subseteq G$
Initialization:
   -If directed is False, duplicate edges in $G$ to include both directions $(i,j)$ and $(j,i)$;
   -Sort edges $G$ by decreasing edge weight $w_{e}$;
   -Iterate through $G$ and construct adjacency list $a$ such that $a[i]=\partial_{i}$. $a[i]$ is sorted by weight since $G$ was sorted;
Algorithm:
-Set $G^{(b)}=[~{}]$;
for $i\leftarrow 1$ to $N$ do
             Compute $\partial_{i}^{(b)}$ by applying Algorithm 1 to $a[i]$;
                 -Ignore initialization step;
                 -Use associated global model ($C/M$) for $\mathcal{L}$;
             Append $\partial_{i}^{(b)}$ to $G^{(b)}$;
end for
   -If directed is False, sort node indices in each edge $(i,j,w_{ij})\in G^{(b)}$ and remove duplicates to convert back to undirected graph;
return $G^{(b)}$;
Algorithm 2 Local backboning (Eq. 10, Eq. 14)
The computational complexity of both Algorithm 1 and Algorithm 2 is $O(E\log E)$, with the bottleneck being the sorting of $G$ during initialization. Both algorithms are therefore only slightly slower in practice than constructing $G$ in the first place. These algorithms are also completely nonparametric, with no required threshold for determining the edges to retain in the backbone, as the optimal edge set is determined automatically through parsimony and the MDL principle.
The final description lengths one is interested in comparing in practice are
  • $\mathcal{L}_{C}^{(\text{global})}(\hat{G}^{(b)})$ (minimum of Eq. II.1)
  • $\mathcal{L}_{C}^{(\text{local})}(\hat{G}^{(b)})$ (minimum of Eq. 10)
  • $\mathcal{L}_{M}^{(\text{global})}(\hat{G}^{(b)})$ (minimum of Eq. 13)
  • $\mathcal{L}_{M}^{(\text{local})}(\hat{G}^{(b)})$ (minimum of Eq. 14)
where $\hat{G}^{(b)}$ is the backbone minimizing the corresponding objective for each expression. The objective giving the lowest description length provides the most parsimonious summary of the data $G$ and should be the preferred backboning model.
It is also useful to construct an inverse compression ratio to determine the extent to which we can compress a network using its MDL-optimal backbones relative to a naïve encoding where we do not use any backbone (equivalent mathematically to setting $G^{(b)}=\{\}$ as the empty graph). For the experiments in Sec. III we will use the microcanonical objectives of Eq. 13 and Eq. 14. (These are, as shown in Appendix A, asymptotically equivalent to the canonical backboning models with geometric weights.) We will compare the levels of compression for the two methods using the inverse compression ratio
$\eta^{(\text{global/local})}$ $=\frac{\mathcal{L}_{M}^{(\text{global/local})}(\hat{G}^{(b)})}{\text{max}[\mathcal{L}_{M}^{(\text{global})}(\{\}),\mathcal{L}_{M}^{(\text{local})}(\{\})]},$ (20)
where global/local indicates the model of interest. Eq. 20 is a useful ratio in practice because:
@@ II.3 Optimization
    It is normalized in $[0,1]$ even for very small networks where $\mathcal{L}_{M}^{(\text{global})}(\{\})$ and $\mathcal{L}_{M}^{(\text{local})}(\{\})$ may differ considerably.
@@
    It is proportional to the final description length values, so that $\eta^{(\text{global})}>\eta^{(\text{local})}$ only when the global method compresses better than the local method in absolute terms (and vice versa for $\eta^{(\text{global})}<\eta^{(\text{local})}$).
Fig. 1 shows a diagram of both the global and local backboning methods for a small synthetic example network, along with the inverse compression ratios for the microcanonical models. Code implementing these algorithms can be found in the PANINIpy package for nonparametric network inference kirkley2024paninipy .
Figure 1: Global and local MDL backbones for a small example network. (a) Schematic of the global backbone objectives (Eq. II.1 and Eq. 13), with the total weight and number of edges indicated for the backbone $G^{(b)}$ (left) and remaining edges (right) of a network $G$. For the canonical objective (Eq. II.1), edge weights $w_{e}$ are sampled according to $w_{e}\sim P(\cdot|\bm{\theta}_{b_{e}})$, where $b_{e}\in\{0,1\}$ is an indicator variable for whether or not $e$ is in the backbone. Meanwhile, for the microcanonical objective (Eq. 14), edge weights are distributed uniformly in the backbone and non-backbone given the constraints imposed by $E^{(b)},W^{(b)}$, which are sampled uniformly given the total weight $W$ and number of edges $E$ in the network. (b) Schematic of the local backbone transmission objectives (Eq. 10 and Eq. 14), applied to out-neighborhoods $\partial_{i}$ for each node $i$. The total strength $s_{i}^{(b)}$ and degree $k_{i}^{(b)}$ are indicated for the backbone neighborhood $\partial^{(b)}_{i}$ around an example node $i$. For both panels, backbone edges are highlighted in blue, while non-backbone edges are highlighted in red. The local network backbone $G^{(b)}$ is the union $G^{(b)}=\cup_{i=1}^{N}\partial^{(b)}_{i}$ of the neighborhood backbones. For this example network, the global method provides a more compressive backbone, giving an inverse compression ratio (Eq. 20) of $\eta^{(\text{global})}=0.71$ versus $\eta^{(\text{local})}=0.89$ for the local backbone.